Find the volume of the region bounded above by the plane z = y/2 and below by the

rectangle.


R ∶ 0 ≤ x ≤ 4,0 ≤ y ≤ 2


1
Expert's answer
2022-03-21T11:35:09-0400

ANSWER:The volume of the region is "V=4"

EXPLANATION

If "V" is the volume of the solid lying vertically above "R" and below the surface "z=f(x,y)" and "f(x,y)\geq0" on "R" , then

"V=\iint_{R}f(x,y) dA"



So , "V=\int_{0}^{2} \int_{0}^{4}\frac{y}{2}dxdy" "V=\int_{0}^{2} \int_{0}^{4}\frac{y}{2}dxdy=\int_{0}^{2}\left ( \frac{y}{2 } \right )\cdot \left ( \int_{0}^{4} dx\right )dy=4\int_{0}^{2}\left ( \frac{y}{2} \right )dy=2\left [ \frac{y^{2}}{2} \right ]_{0}^{2}=\left ( 2^{2} \right )-0=4"